OpenStudy (anonymous):

cos x/2 = cos x. find all solutions in the interval [0,2π)?

OpenStudy (anonymous):

The general solution for cos(x)=a, is \[n \pi \pm a\] Therefore, \[x/2 = n \pi \pm x\] \[x = 2n \pi \pm 2x\] \[3x = 2n \pi | -x = 2n\pi\] \[x = 2/3 n \pi | x = -2n \pi\] \[[0,2 \pi], x=0 , 2 \pi / 3 , 4 \pi/3 \] This is the solution set!

OpenStudy (anonymous):

Thank you so much!